OpenStudy (anonymous):

The two angles have equal measures so 3x + 38 = 5x + 20. Solve for x and use your result to find the measures of the angles. You need to plug in 9 to find the measurement of the angles in degrees. Can somebody help me with this problem.?

OpenStudy (anonymous):

they told you what x equals. x=9. The measure of the angle is 3(9)+38=5(9)+20=65º

OpenStudy (anonymous):

Thank u so much!